=== Current encoding is LO1 (LO1).

— Simple Tests — === Test using LO1 font encoding, accent macros, Latin-1 chars: täter élève élève === Test using LO1 font encoding, german hyphenation patterns: täter élève === Test using LO1 font encoding, french hyphenation patterns: täter élève === Test using OT1 font encoding, french hyphenation patterns: täter élève délégué — Major Tests —

— T1: À Á Â Ã Ä Ă Å à á â ã ä ă å Ç Ć Č ç ć č Ď Ğ ğ È É Ê Ẽ Ë Ě è é ê ẽ ë ě Ì Í Î Ĩ Ï İ ì í î ĩ ï i ì í î ĩ ï i̇ Ñ Ń Ň ñ ń ň Ò Ó Ô Õ Ö Ő ò ó ô õ ö ő Ŕ Ř ŕ ř Ś Š Ş ś š ş Ť Ţ ţ Ù Ú Û Ũ Ü Ű Ů ù ú û ũ ü ű ů Ý ý Ź Ž Ż ź ž ż composite glyphs: b̈ B̈ b́ B́ ḃ Ḃ b̄ B̄ b̂ B̂ b̀ B̀ b̃ B̃ b̋ B̋ b̆ B̆ b̌ B̌ b̊ B̊ b̄ B̄ b̧ B̧ b. B. special glyphs: ÆŒØæıȷœøß—–¡¿“”‘’Łł$£

— OT1: À Á Â Ã Ä Ă Å à á â ã ä ă å Ç Ć Č ç ć č Ď Ğ ğ È É Ê Ẽ Ë Ě è é ê ẽ ë ě Ì Í Î Ĩ Ï İ ì í î ĩ ï i̇ ì í î ĩ ï i̇ Ñ Ń Ň ñ ń ň Ò Ó Ô Õ Ö Ő ò ó ô õ ö ő Ŕ Ř ŕ ř Ś Š Ş ś š ş Ť Ţ ţ Ù Ú Û Ũ Ü Ű Ů ù ú û ũ ü ű ů Ý ý Ź Ž Ż ź ž ż composite glyphs: b̈ B̈ b́ B́ ḃ Ḃ b̄ B̄ b̂ B̂ b̀ B̀ b̃ B̃ b̋ B̋ b̆ B̆ b̌ B̌ b̊ B̊ b̄ B̄ b̧ B̧ b. B. special glyphs: ÆŒØæıȷœøß—–¡¿“”‘’Łł$£

— LO1: À Á Â Ã Ä Ă Å à á â ã ä ă å Ç Ć Č ç ć č Ď Ğ ğ È É Ê Ẽ Ë Ě è é ê ẽ ë ě Ì Í Î Ĩ Ï İ ì í î ĩ ï i̇ ì í î ĩ ï i̇ Ñ Ń Ň ñ ń ň Ò Ó Ô Õ Ö Ő ò ó ô õ ö ő Ŕ Ř ŕ ř Ś Š Ş ś š ş Ť Ţ ţ Ù Ú Û Ũ Ü Ű Ů ù ú û ũ ü ű ů Ý ý Ź Ž Ż ź ž ż composite glyphs: b̈ B̈ b́ B́ ḃ Ḃ b̄ B̄ b̂ B̂ b̀ B̀ b̃ B̃ b̋ B̋ b̆ B̆ b̌ B̌ b̊ B̊ b̄ B̄ b̧ B̧ b. B. special glyphs: ÆŒØæıȷœøß—–¡¿“”‘’Łł$£

— End of Tests —